OpenStudy (anonymous):

PLEASE HELP CALCULUS: Use the Generalized Power Rule to find the derivative of the function. y=[1/(w^3-1)]^8

OpenStudy (anonymous):

\[y=\left( \frac{ 1 }{ w^3-1 } \right) ^8\]

OpenStudy (nincompoop):

okay you need quotient and chain rule

OpenStudy (nincompoop):

rules*

OpenStudy (anonymous):

Or you can make it to the power of -1 and use chain rule/ power rule st00f, as it suggests.

OpenStudy (nincompoop):

\[ y = \frac{1^8}{(w^3-1)^8} \]

OpenStudy (nincompoop):

perform quotient rule and power rule

OpenStudy (anonymous):

so this was my answer \[y=8(w ^{-3}-1)^7(-3w ^{-4})\]

OpenStudy (nincompoop):

I don't really care about the answer that you said was already wrong

OpenStudy (anonymous):

Your derivative is wrong, just do what nin asks D: he's fat

OpenStudy (anonymous):

ok, I dont understand what u want me to with that?

OpenStudy (nincompoop):

\[ u = 1^8 \\ u' = 0 \\ v = (w^3-1)^8 \\ v' = 8(w^3-1)^7 \times 2w^2 \\ y' = \frac{u'v - uv'}{v^2} \]

OpenStudy (nincompoop):

sorry

OpenStudy (anonymous):

oh the quotient rule. I forgot about that.

OpenStudy (nincompoop):

\[ v' = 8(w^3-1)^7 \times 3w^2 \]

OpenStudy (nincompoop):

went trigger happy with the 2 there sorry for the typo

OpenStudy (anonymous):

ok the reason I didn't approach with the quotient rule was because with the demonstration of webassign, it was showing me to approach the equation by y=n(blah blah)^n-1(dy/dx)